OpenStudy (anonymous):

Solve the following system: y = x + 3 2x + y = 9 (2, 5) (5, 2) (-2, 5) (2, -5)

OpenStudy (anonymous):

i think its a

OpenStudy (matlee):

x+3=9-2x +x +x 3=9-1x -9 -9 -12=-1x x=12 So you plug it in and the right answer is -2,5

OpenStudy (studygurl14):

(-2,5) doesn't work...

OpenStudy (studygurl14):

\(\large (5)\neq (-2)+3\)

OpenStudy (anonymous):

i think its a right?

OpenStudy (matlee):

OK fine its B

OpenStudy (matlee):

i mean D

OpenStudy (studygurl14):

No, @stupid12 is correct. It's A

OpenStudy (anonymous):

hey if you go throught and plug all the answer choices in you should find the correct answer

OpenStudy (anonymous):

here is how you do it though

OpenStudy (anonymous):

yeah i think it is because its the only one that works in the equation

OpenStudy (studygurl14):

Good job @stupid12 @matlee you added x to both sides of the equation instead of subtracting.

OpenStudy (anonymous):